# SimMLM: A Simple Framework for Multi-modal Learning with Missing Modality

### v1

@@ 3.3 Dynamic Mixture of Modality Expert (DMoME)
Notably, the input to the gating network can be an incomplete set of modalities. If certain modalities are missing, those inputs are treated as $\mathbf{0}$, and the gating values for the corresponding missing modality experts’ outputs are set to $\infty$- ∞ (which causes the corresponding weight values to be 0 after the softmax). Mathematically, the gating outputs are set as:
$g_{m}=\begin{cases}G(\mathbf{x};\phi)^{(m)},&\text{if modality }m\text{ is present}\\ -\infty,&\text{if modality }m\text{ is missing}\end{cases},$ (1)
@@ 4.1 Datasets
avMNIST (Audiovision MNIST [34]) is a digit classification dataset comprising 1,500 image-audio pairs. The visual modality consists of $\times 28$ 28 grayscale images of handwritten digits (0-9) from MNIST [12], while the audio modality is sourced from the Free Spoken Digits Dataset***Dataset link: https://github.com/Jakobovski/free-spoken-digit-dataset..
@@ 4.2 Implementation Details
BraTS 2018: We employed nnUNet [8], a widely used and powerful medical image segmentation backbone, as the expert network for each modality, which can be viewed as a 3D U-Net. The gating network $G$ in DMoME utilized lightweight CNNs to capture broad contextual information, followed by a linear layer generating 4×34$\times$ 34 × 3 weights for the four modalities and three segmentation tasks. We optimized both training stages using the Adam optimizer [11] ($lr=0.01$). The $\mathcal{L}_{task}$ in Eq.4 combined Dice [23] and binary cross-entropy losses.
@@
avMNIST: Following [20, 35], we adopted a LeNet-5 [12] backbone for both image and audio modalities, with audio converted into $\times 20$ 20 MFCC spectrograms [33]. The gating network $G$ followed a similar design as in BraTS 2018. Two-stage training was conducted using Adam ($lr=0.001$) with cross-entropy loss as $\mathcal{L}_{task}$.
@@ 5 Analysis
Table 3: Comparison of model accuracy (Dice score $\uparrow$, normalized to 100%, averaged across all missing & full modality settings) and efficiency (#\## of parameters and FLOPS) on BraTS 2018 official-evaluated benchmark.

### v2

@@ 3.3 Dynamic Mixture of Modality Expert (DMoME)
Notably, the input to the gating network can be an incomplete set of modalities. If certain modalities are missing, those inputs are treated as $\mathbf{0}$, and the gating values for the corresponding missing modality experts’ outputs are set to -$\infty$ (which causes the corresponding weight values to be 0 after the softmax). Mathematically, the gating outputs are set as:
[…]
@@ 4.1 Datasets
avMNIST (Audiovision MNIST [35]) is a digit classification dataset comprising 1,500 image-audio pairs. The visual modality consists of $28\times 28$ grayscale images of handwritten digits (0-9) from MNIST [13], while the audio modality is sourced from the Free Spoken Digits Dataset***Dataset link: https://github.com/Jakobovski/free-spoken-digit-dataset..
@@ 4.2 Implementation Details
BraTS 2018: We employed nnUNet [9], a widely used and powerful medical image segmentation backbone, as the expert network for each modality, which can be viewed as a 3D U-Net. The gating network $G$ in DMoME utilized lightweight CNNs to capture broad contextual information, followed by a linear layer generating $4\times 3$ weights for the four modalities and three segmentation tasks. We optimized both training stages using the Adam optimizer [12] ($lr=0.01$). The $\mathcal{L}_{task}$ in Eq.4 combined Dice [24] and binary cross-entropy losses.
@@
avMNIST: Following [21, 36], we adopted a LeNet-5 [13] backbone for both image and audio modalities, with audio converted into $20\times 20$ MFCC spectrograms [34]. The gating network $G$ followed a similar design as in BraTS 2018. Two-stage training was conducted using Adam ($lr=0.001$) with cross-entropy loss as $\mathcal{L}_{task}$.
@@ 5 Analysis
Table 3: Comparison of model accuracy (Dice score $\uparrow$, normalized to 100%, averaged across all missing & full modality settings) and efficiency (#\# of parameters and FLOPS) on BraTS 2018 official-evaluated benchmark.
@@ Acknowledgements
Supplementary Material
Implementation of the gating network
The gating network $G$ in our DMoME framework is designed with a generic and efficient architecture to ensure adaptability across diverse tasks and modalities. At a high level, it consists of a feature extraction block (e.g., 2D or 3D convolutional neural networks (CNNs), transformers, etc.), followed by a linear layer that generates low-dimensional weight vectors. This architecture strikes a balance between simplicity and flexibility, enabling it to adjust modality contributions effectively across tasks. Furthermore, its modular design allows seamless accommodation of varying task requirements and diverse input modalities, ensuring broad applicability and adaptability.
As shown in Table A6, for the BraTS 2018 segmentation task, the input modalities are four input imaging modalities ($4\times 128\times 128\times 128$ with zeros if certain modalities are missing). Here, since they are all images sharing the same dimensionality, we simply extract 256dim256-dim features using a 5-layer 3D CNN to directly fuse visual information and send the features to an linear layer to get a $4\times 3$ gating vector (# modalities $\times$ #tasks). The vector will be transformed using the softmax function to generate gating weights, which adjust the contribution of each modality per task. In contrast, the avMNIST classification task involves heterogeneous modalities: the image modality ($1\times 28\times 28$) and the audio modality ($1\times 20\times 20$). To account for this heterogeneity, two separate 2-layer CNNs are used to extract 128-dimensional feature vectors for each modality. These features are concatenated to form the same size of 256dim256-dim representation, which is then processed by a similar linear layer to generate a $2\times 1$ gating vector. This modular approach ensures flexibility and robust feature integration across diverse data types.
In this work, we simply designed these two lightweight gating networks empirically, which provides substantial improvements. Thanks to the flexibility of the framework, readers are encouraged to explore more advanced architectures for their customized tasks.
General design BraTS 2018 avMNIST
Feature extractor
4-modality MRIs
$4\times 128^{3}$
image
$1\times 28^{2}$
audio
$1\times 20^{2}$
5-layer
CNNs (3D)
2-layer
CNNs (2D)
2-layer
CNNs (2D)
Flatten Flatten & concat
256-dim feature 256-dim feature
Linear layer
Linear layer
Output size: $4\times 3$
(4 modalities $\times$ 3 tasks)
Linear layer
Output size: 2 $\times$ 1
(2 modalities $\times$ 1 task)
Table A6: Designs of the gating network $G$ for the BraTS 2018 segmentation task and the avMNIST classification task.
Importance of the design choice and the training recipe in SimMLM
Dice score $\uparrow$
Method ET TC WT
DMoME w/o expert pretraining 62.24 77.42 86.63
DMoME w/ static averaging 61.40 77.49 86.70
DMoME w/ dynamic weighting (SimMLM) 63.22 78.54 87.21
Table A7: Importance of the design choice and the training recipe in SimMLM. Reported values are the average segmentation performance on the BraTS 2018 validation set, across all missing & full modality settings.
In this section, we would like to highlight the contribution of pretraining and dynamic weighting by comparing our proposed DMoME with its two variants:
  • DMoME w/o expert pretraining: Here, all modality experts and the gating network $G$ were directly co-trained from random initialization, without pretraining. The total number of training epochs is the same as the proposed two-stage one for fair comparison.
  • DMoME w/ static averaging: In this setting, the gating network $G$ is replaced with a simple static averaging operation. The final output can be simply viewed as the mean of the available modality experts’ outputs.
Results in Table A7 demonstrate that both expert pretraining and dynamic reweighting contribute to the success of SimMLM. Skipping expert pretraining (DMoME w/o expert pretraining) results in a noticeable performance drop, highlighting the critical role of independent pretraining. This step enables modality experts to focus on task-relevant knowledge without interference from cross-expert interactions, which could be noisy especially at the beginning of training. Replacing the dynamic gating mechanism with static averaging (DMoME w/ static averaging) also causes a significant performance drop across all subregions (ET, TC, WT). This underscores the effectiveness of dynamically assigning weights to modality experts, which is not just a simple model ensembling strategy.
Impact of performing dynamic weighting at different levels in DMoME
As shown in Figure A5, we propose a simple yet effective mixture of modality experts strategy, which directly applies expert weighting at the logit level (before softmax) instead of probabilities.
Figure A5: Three types of weighted sum micro design: feature-level, logit-level (SimMLM), and probability-level.
Here, we tested the importance of this design choice, by comparing it to the other two alternatives with the same gating network design:
  • Feature-level mixture: Here, the dynamic weighting is applied at hidden feature maps (before passing through the segmentation/classification head) rather than the output space, which is adopted in MoMKE [39]. The feature maps from modality experts are later mixed with the weights produced by the gating network. A segmentation head is followed to generate the final results given the mixed feature map.
  • Probability-level mixture: Here the dynamic weighting is applied to the probabilities from each expert (after softmax).
Results in Table A8 show that performing logit-level mixture generally achieves better results, especially on the more challenging enhancing tumor (ET) and tumor core (TC) segmentation tasks. We believe this improvement stems from SimMLM’s logit-level weighting, where the weighting parameter functions as a temperature to rescale the confidence of each modality’s prediction in the final output. This approach aligns with the principles of temperature rescaling [6] for model calibration, which can effectively adapt model predictions to better reflect the likelihood of ground truth correctness, leading to more reliable multi-modal inference especially in challenging scenarios. In our main paper, we have already demonstrated that our DMoME framework, when used standalone, can achieve much lower calibration errors compared to MoMKE in Table 5, thanks to the logit-level dynamic weighting.
Dice score $\uparrow$
Method ET TC WT
Feature-level mixture 62.76 78.19 87.27
Probability-level mixture 62.72 78.31 87.33
Logit-level mixture (SimMLM) 63.22 78.54 87.21
Table A8: Importance of performing dynamic weighting at logit level. Reported values are the average segmentation performance on the BraTS 2018 validation set, across all missing & full modality settings.
Figure A6: Model performance comparison with varying rates of missing audio data during training on av-MNIST.
Figure A7: MoFe weight $\lambda$ tuning on avMNIST, UPMC Food-101.
SimMLM’s robustness against varied levels of missing modalities at training time
To further investigate SimMLM’s resilience to training-time missingness, we conducted experiments where the audio coverage rate was reduced to 5%, 10%, 20%, 30%, 50% during the training. We then evaluated the model on the full modality evaluation set and compared its performance to ShaSpec [36] and MoMKE [39]. As shown in Figure A6, our model consistently achieves the highest accuracy in all settings, demonstrating substantial improvement in handling training-time modality absence. Remarkably, our model achieves even better accuracy comparable to competing methods that use 50% of the audio data, with only 30% of it during training (see red dashed line). These results demonstrate our method’s potential to handle training-time modality missingness as well.
MoFe coefficient tuning on avMNIST and UPMC Food-101
Setting the MoFe loss coefficient to 0.1 generalizes well across tasks. As shown in Fig. A7, where MoFe consistently improves performance across a range of values, with 0.1 yielding stable results.
Discussion: How well does the unimodal performance align with the importance derived from the learned gating weights?
Interestingly, the gating weights show weak to moderate alignment with unimodal performance at the sample level. On UPMC Food-101, the Pearson correlation between each modality’s gating weight and its unimodal confidence on the correct class is 0.29 for image and 0.41 for text. This partial alignment suggests that while unimodal confidence reflects individual modality strength, it may not fully capture cross-modal complementarity or sample difficulty—both of which influence the model’s reliance during inference. We consider this behavior desirable, as it indicates the gating mechanism is context-aware rather than simply mirroring unimodal confidence.
Clinical value of SimMLM
In real-world clinical scenarios, especially in low-resource settings, certain imaging modalities may be unavailable due to technical issues, equipment limitations, or patient factors. SimMLM’s ability to generate more reliable and accurate segmentation under these conditions is essential. The model not only provides direct importance values for each input modality but also generates more informative confidence maps, including entropy maps, for voxel-wise predictions in the output space. This capability is particularly valuable in critical applications like tumor detection and diagnosis, where accurate and detailed segmentation is crucial for treatment planning and patient care. By enabling clinicians to make confident decisions—whether by providing insights or suggesting further investigation or special care for uncertain regions—even in the presence of incomplete data, SimMLM greatly enhances the robustness and reliability of clinical workflows. In the future, we will explore SimMLM’s application in broader, high-stakes scenarios, such as surgical scene reconstruction, action segmentation, and surgical planning.